Tom is 4 years older than Jerry. In 2 years Jerry will be half as old as Tom. Find their ages now.
Answer by Cromlix(4381)

Hi there,
Tom is 4 years older than Jerry.
T = 4 + J
or
J = T - 4
In 2 years Jerry will be half as old as Tom.
J + 2 = 1/2(T + 2)
Replace J with T - 1
T - 4 + 2 = 1/2T + 1
Collect like terms
T - 1/2T = -2 + 4 + 1
1/2T = 3
T = 6
Tom = 6 years old
Jerry = 2 years old.
Hope this helps :-)